can the sum of the magnitude of two vector ever bi equal to the magnitude of the sum of two vector

Grade:12th pass

1 Answers

Arun
5 years ago
If we alligntwo vectors in same direction we can get “the sum of magnitude of two vector ever bi equal to the magnitude of sum of these vector“for example A = i, B= 2i|A|+|B|=1+2 = 3|A+B| = |i+2i| = 3  RegardsArun (askIITians forum expert)